# ProveRAG: Provenance-Driven Vulnerability Analysis with Automated Retrieval-Augmented LLMs

### v1

@@ 4. ProveRAG Methodology
ProveRAG works based on two main stages: The generation module and the evaluation module. Figure 1 shows the overall architecture. First, in the generation module, the Retr. LLM will load web content from external sources (starting from NVD) about a CVE vulnerability, and it will summarize the content with respect to the exploitation and mitigation information. Next, the Gen. LLM will look at the summaries and generates detailed exploitation and mitigation strategies for that specific CVE. Furthermore, in the evaluation module, the Eval. LLM will critique its own response by looking at the content from the official and external (NVD, CWE, and hyperlinks in NVD) websites to show evidence on whether its response were accurate, hallucinated, or omitted critical information. Finally, the LLM will provide provenance by showing similar pieces of information from both the response and the context while providing rationale to ensure trust. It is worth noting that we used the OpenAI’s gpt-4o-mini (OpenAI, 2024) model as the LLM for ProveRAG. This model’s training data knowledge is up to Oct. 2023, with 128,000 tokens for context window (OpenAI, 2024).
4.1. System Design
4.1.1. Generation Module:
@@
This module is to query the Gen. LLM to provide exploitation and mitigation responses for a CVE (refer to Figure 1). The following is the generation prompt:
@@
We test on two retrieval techniques (to address RQ2). Figure 2 shows how each of these techniques is used. For the chunking technique, the LLM will use top-10 most similar chunks of 15,000 characters from the resources. Second, for the summarizing technique, the Retr. LLM goes through the NVD, then to the CWE and hyperlinks to summarize the content for the exploitation and mitigation questions (see Appendix 5.2.1 for the full prompts).
@@
In the chunking technique, we split the content of all the URLs into smaller chunks and embed these chunks as vectors using OpenAI’s text-embedding-ada-002 embedding model (by utilizing the LangChain framework (LangChain, 2024)). These embeddings are then indexed to facilitate efficient retrieval. During runtime, the user’s query is also vectorized using the same embedding model, and a similarity search is conducted against the indexed chunks. The top-10 results are retrieved and fed as context into the Gen. LLM’s prompt, allowing the model to answer the user’s query with the most relevant content. However, this approach requires a very large context window and can include redundant information that might mislead the LLM.
@@
Therefore, we design an automatic summarization technique, inspired by (Kim et al., 2024; Jiang et al., 2024), for a specific CVE (which will be further integrated into ProveRAG’s architecture). We start by exploring the NVD website and summarize the content with respect to a specific CVE-ID, then move to the CWE link, and finally to the hyperlinks (available in NVD) to provide a summary with respect to the exploitation and mitigation questions (see Appendix A for the specific summarization prompts used). A key advantage of the summarizing technique is its efficiency in handling large-scale data without relying extensively on document chunking, which enhances retrieval quality and accuracy of responses. This makes the system particularly suited for addressing more complex cybersecurity queries, such as those related to CVE mitigation that require exploration of more external resources.
Figure 2. Chunking vs. Summarizing for retrieval process used in both generation and evaluation of ProveRAG.
4.1.2. Evaluation Module:
@@
This module is for the LLM to self-critique its own response (to answer RQ3). We carefully and iteratively designed structured, detailed, and succinct prompts with the chain-of-thought technique (Wei et al., 2022) to guide the model to think before answering while providing evidence. The following is the evaluation prompt:
@@
Evaluation Prompt For the [exploitation/mitigation] information of [CVE-xxx-xxx]: Please evaluate the response based on the provided evidence and select the appropriate value for the following attributes: ————————- response: [LLM response] ————————- evidence: [content from NVD + CWE + Relevant Refs.)] ————————-
@@
There are three attributes that the LLM is tasked to provide a response, and they are: value, rationale, and provenance prompts. The exact prompts provided to the LLM are shown in Appendix B. For these attributes, we ask the LLM to provide a value: True Positive (TP) - indicating accurate prediction, False Positive (FP) - indicating the detection of hallucination, or False Negative (FN) - indicating failed retrieval given the response and the evidence (from NVD+CWE+hyperlinks). Next, the LLM is asked to provide a rationale for the selected value, and finally the provenance attribute where the LLM is tasked to show evidence for its response.
@@
5. Evaluation of ProveRAG
To answer each research question, we discuss the curation of the dataset for (RQ1) and provide results for different RAG techniques (RQ1 and RQ2). Furthermore, we support the efficacy of provenance by using evaluation metrics (RQ3), show how each URL attributed to the Gen. LLM’s response (RQ3), and finally we will show some specific examples of ProveRAG’s responses.
Table 1. Comparison between Summarizing and Chunking techniques for retrieving the content for CVEs (Provenance Features: TP: accurate prediction, FP: detected hallucination, FN: failed retrieval)
@@ 5. Evaluation of ProveRAG
5.1. Dataset & Experiment Design
@@
5.2. Effectiveness of ProveRAG
5.2.1. Chunking vs. Summarizing Techniques
@@ 5.2. Effectiveness of ProveRAG
From the results, first we can see that for the prompt-only experiment, the performance is poor, for both exploitation and mitigation results. This shows the big limitation of LLMs’ knowledge with new, unseen data, as the entire dataset consists of 2024 vulnerabilities. This goes back to RQ1 and how we can deal with this problem. By integrating the retrieval techniques, we can see that the temporal alignment of the LLM is improved by the inclusion of the curated dataset, indicating that such an approach may alleviate limitations related to temporal alignment for cybersecurity applications. For the exploitation results, all evaluated methodologies exhibited high accuracy, especially for NVD-Only which was limited to the information about the CVE’s exploitation on the NVD website. Another observation is that adding more data slightly decreases the chunking technique’s performance. Conversely, for the more complex queries (i.e., mitigation information), significant variations are observed. The NVD-Only configuration achieves around 29% accuracy in providing mitigation strategies, which makes sense given that the NVD website lacks a dedicated mitigation section; yet, it does offer some recommendations for certain CVEs. When the CWE is added into the generation of results, we can see a big increase in TPs (from 139 to 262). However, adding additional URLs into the generation mix appears to diminish performance by about 7%, suggesting that an excess of links might degrade the quality of the retrieved data segments with the chunking technique. This indicates that providing excessive context can degrade performance when chunking is used.
In contrast, the summarizing technique significantly outperforms the chunking technique when provided with the same set of links for data generation, showing an increase from 227 to 379 in TPs — a rise of over 30%. Therefore, these findings strongly support the conclusion that summarization is a better and more efficient technique. As a result, it has been incorporated into ProveRAG to help align the system to vulnerabilities discovered after training cut-off. Furthermore, this technique alleviates the context window problem (RQ2) for vulnerability analysis, as only the essential information (i.e., summaries) are put into the Gen. LLM’s prompt.
5.2.2. ProveRAG Results
@@
Table 2. ProveRAG results for the 482 CVEs
@@
5.3. Evaluation Metrics for Provenance
@@
5.3.1. Attribution:
To assess the attribution of each URL summary (i.e., NVD, CWE, and hyperlinks) and how much each contributed to the generation of the results, we made use of the Captum library (Miglani et al., 2023). This library offers implementations of various perturbation- and gradient-based attribution algorithms for PyTorch models. Attribution methods allow for model explainability by assigning scores to each input feature, representing its impact on the model’s output. In our experiment, we compute attribution scores using feature ablation, a perturbation-based method that evaluates the model on altered inputs. This method replaces each feature with its corresponding value from the baseline feature vector to assess its influence. Equation 1 assesses the contribution of different features to the final result.
(1) $\Delta y=y(x)-y(x_{\setminus F})$
@@ 5.3.1. Attribution:
Figure 3. ProveRAG Exploitation & Mitigation Cosine Similarity and Rouge-L Values for #TPs, #FPs, and #FNs
5.3.2. Rouge-L and Cosine Similarity:
@@
To evaluate the quality of the Provenance results, we incorporated Rouge-L and Cosine Similarity (based on embeddings) metrics into our analysis. Rouge-L focuses on the longest common subsequence, which measures the sequence similarity between the generated responses and the reference texts. On the other hand, Cosine Similarity leverages a Sentence Transformer model, specifically the multi-qa-mpnet-base-dot-v1, which is optimized for semantic search (SentenceTransformers, 2024). This model utilizes a 512-context window to compare the semantic congruence between the response and the evidence part of the provenance, ensuring that the evaluation captures not just the textual, but also the contextual and semantic accuracy of the information provided.
5.4. Provenance Quality
@@
To answer RQ3, we applied Rouge-L and Cosine Similarity to measure the quality of both the provided responses and their supporting provenance. These metrics help us assess the alignment between the system’s outputs and verifiable evidence when evaluating TPs, FPs, and FNs. Figure 3 illustrates these values. As anticipated, TPs consistently achieve the highest scores across both metrics, indicating greatest overlap between the generated response and the supporting evidence, which underscores the effectiveness of ProveRAG. Note that the scores of FNs is close to that of TPs. This is expected because the FNs often just omitted some details in the response when comparing it with the source context. There is no value for FN in the exploitation part, as the LLM made no FN predictions in that category. Detailed examples are provided in Section 5.6 and in Appendix C to contextualize these values.
@@
Furthermore, to analyze the effect of summary for each URL to the LLM’s response, we made use of the Captum library (Miglani et al., 2023). We calculate attribution by using the summaries from each relevant website as context for the generation prompt. If a website is deemed irrelevant, its attribution score is set to zero, as it does not provide usable information for the LLM. To facilitate comparison across CVEs, the attribution scores are normalized to unit norm. Figure 4 shows these attribution values. From the results, it is evident that for exploitation, the majority of information in the summaries originates from NVD. For mitigation summaries, most of CWE’s attention scores are higher than 0.9, whereas the values for NVD and the hyperlinks stay mostly unchanged. This indicates that the Gen. LLM put the most attention to the NVD summary for generating the exploitation response, and to the CWE summary for generating the mitigation strategies.
Figure 4. ProveRAG Attribution Results
5.5. Improvement w/ Better Structured Context
@@
In Sec. 5.2.2, we demonstrated that ProveRAG achieves 95% and 80% TPs (with NVD, CWE, and all hyperlinks in the NVD website). Therefore, we raise the question on whether this can be further improved. Note that CWE is addressing mitigation at the weakness level, not the vulnerability level. Through our research, we found Aqua Vulnerability Database (Security, 2024), which gives a highly structured and detailed context for each CVE’s exploitation steps and mitigation strategies. Thus, we experimented ProveRAG by purposefully including Aqua in the retrieval summarizing process, and into the evaluation module for provenance. The results are shown in the bottom row of Table 2. The results demonstrate the significant increase in the exploitation and mitigation detection of ProveRAG-Aqua, with 99% and 97% for TPs, respectively. This shows that the Aqua database plays a pivotal role in enhancing ProveRAG’s performance by providing more detailed and structured context. These results are significantly better than the baseline experiments, where prompt-only achieved just 6% accuracy for mitigation strategies, and the retrieval chunking technique reached about 47%.
Table 3. Critical CVEs Relevancy with respect to the URLs
@@
Figure 5. ProveRAG-Aqua attribution results
@@
To further evaluate Aqua’s impact on the Gen. LLM’s responses, Figure 5 presents the attribution of each summary in ProveRAG-Aqua. The results reveal that, for the mitigation prompt, both Aqua and CWE exhibit a wide distribution of attention scores, with Aqua having a higher median around 0.7 compared to CWE’s median of 0.43. For the exploitation prompt, both Aqua and NVD display a broad range of attention scores, mostly positive, with Aqua again leading with a median near 0.7, while NVD’s median is around 0.3. This shows that when Aqua is included, it is the dominant source for the Gen. LLM for both exploitation and mitigation questions, with some contribution from NVD for exploitation and from CWE for mitigation strategies.
5.6. Specific Examples
In this section, we show how the evaluation module of ProveRAG works (RQ3). Additional examples are provided in Appendix C as well. Here is one example that shows a TP response:

### v2

@@ 4 ProveRAG Methodology
ProveRAG works based on two main stages: The generation module and the evaluation module. Figure 1 shows the proposed architecture. Note that the Retr. LLM, Gen. LLM, and Eval. LLM are the same model. First, in the generation module, the Retr. LLM will load web content from external sources (starting from NVD) about a CVE vulnerability, and it will summarize the content with respect to the exploitation and mitigation information. Next, the Gen. LLM will look at the summaries and generates detailed exploitation and mitigation strategies for that specific CVE. Furthermore, in the evaluation module, the Eval. LLM will critique its own response by looking at the content from the official and external (NVD, CWE, and hyperlinks in NVD) websites to show evidence on whether its response were accurate, hallucinated, or omitted critical information. Finally, the LLM will provide provenance by showing similar pieces of information from both the response and the context while providing rationale to ensure trust. It is worth noting that we used the OpenAI’s gpt-4o-mini [25] model as the LLM for ProveRAG. This model’s training data knowledge is up to Oct. 2023, with 128,000 tokens for context window [25]. However, to show the proposed ProveRAG framework can be generalized across other LLMs, we compare the performance of gpt-4o-mini with Llama-3.1-8B [17] in Sec. 5.6. The Llama-3.1-8B model is selected as it has the same context window of 128,000 tokens, and knowledge cutoff until Dec. 2023 (no exposure to CVEs from 2024).
4.1 System Design
4.1.1 Generation Module:
@@
This module is to query the Gen. LLM to provide exploitation and mitigation responses for a given CVE (refer to Figure 1). The following is the generation prompt:
@@
We test on two retrieval techniques (to address RQ2). Figure 2 shows how each of these techniques is used. For the chunking technique, the LLM will use top-10 most similar chunks of 15,000 characters from the resources. Second, for the summarizing technique, the Retr. LLM goes through the NVD, then to the CWE and hyperlinks to summarize the content for the exploitation and mitigation questions The following shows the prompt used in Retr. LLM to assess each URL’s relevancy, and, if relevant, the LLM will summarize the URL content specifically to address either the Exploitation or the Mitigation question.
Relevancy Prompts Exploitation Prompt You are a cybersecurity expert. Your task is to analyze the provided URL content for the [CVE-xxx-xxx] and provide a detailed summary. Content: content Please follow the steps below: Step 1: Assess Relevancy - Does the content provide relevant information to describe how this [CVE-xxx-xxx] can be exploited? - Answer: [Yes/No] Step 2: Summarize Relevant Information - If the answer is "Yes" in Step 1: - Summarize the content with step-by-step description to exploit this vulnerability. - If the answer is "No" in Step 1: - Summary: NONE Mitigation Prompt You are a cybersecurity expert. Your task is to analyze the provided URL content for the [CVE-xxx-xxx] and provide a detailed summary. Content: content Please follow the steps below: Step 1: Assess Relevancy - Does the content provide relevant information to describe the potential mitigation strategies for this vulnerability? - Answer: [Yes/No] Step 2: Summarize Relevant Information - If the answer is "Yes" in Step 1: - Summarize the content to describe the potential mitigation strategies for this CVE. - If the answer is "No" in Step 1: - Summary: NONE
@@
In the chunking technique, we split the content of all the URLs into smaller chunks and embed these chunks as vectors, and these embeddings are then indexed to facilitate efficient retrieval. We further show the performance for different models in Sec. 5.6. During runtime, the user’s query is first vectorized, and a similarity search is conducted against the indexed chunks. The top-10 results are retrieved and fed as context into the Gen. LLM’s prompt, which allows the model to answer the user’s query with the most relevant content. However, this approach requires a very large context window and can include redundant information that might mislead the LLM.
@@
Therefore, we design an automatic summarization technique, inspired by [14, 10], for a specific CVE (which is integrated into ProveRAG’s architecture). We start by exploring the NVD website and summarize the content with respect to a specific CVE-ID, then move to the associated CWE link, and finally to the hyperlinks (available in NVD) to provide a summary with respect to the exploitation and mitigation questions. A key advantage of the summarizing technique is its efficiency in handling large-scale data without relying extensively on document chunking, which enhances retrieval quality and accuracy of responses. This makes the system particularly suited for addressing more complex cybersecurity queries, such as those related to CVE mitigation that require exploration of more external resources.
4.1.2 Evaluation Module:
@@
This module is for the LLM to self-critique its own response (to answer RQ3). We carefully and iteratively designed structured and succinct prompts with the chain-of-thought technique [35] to guide the model to think before answering while providing evidence. The following is the evaluation prompt:
@@
Evaluation Prompt For the [exploitation/mitigation] information of [CVE-xxx-xxx]: Please evaluate the response based on the provided evidence and select the appropriate value for the defined attributes: ————————- response: [LLM response] ————————- evidence: [content from NVD + CWE + Relevant Refs.)] ————————-
The LLM is tasked with providing responses for three attributes: value, rationale, and provenance. The following are these attributes for the Eval. LLM:
Evaluation Module Value prompt - Description: This attribute represents the accuracy of the response based on the provided context about the CVE-ID. - Allowed Values: - ‘TP’ (True Positive): The response fully and accurately reflects the information in the context. - ‘FP’ (False Positive): The response contains inaccurate or incorrect information that are not supported by the context. - ‘FN’ (False Negative): The response omits information that is present in the context. - Requirements: - Must strictly be one of the allowed values: ‘TP’, ‘FP’, or ‘FN’. - The selection must be based on an objective comparison between the response and the context. - Must follow the guidelines for Provenance. Rationale prompt - Description: This attribute provides the reasoning behind the selected "value" for the evaluation of the response.
Evaluation Module Provenance prompt - Detailed Instructions for Provenance based on the selected "value": - If "value" is ‘TP’: 1. Carefully compare the response with the context. 2. Identify and extract the key segments from both the response and the context that align perfectly. 3. Format the provenance as follows: - ‘response: "[Extracted segment from the response that matches the context]".’ - ‘context: "[Corresponding segment(s) from the context]".’ - If "value" is ‘FP’: 1. Review the response thoroughly to identify parts that are incorrect or not present in the context. 2. Format the provenance as follows: - ‘response: "[Incorrect or unsupported segment from the response]".’ - ‘context: "[The closest matching part from the context, or indicate ‘No corresponding information in context’ if none exists]".’ - If "value" is ‘FN’: 1. Examine the context to find essential information that is present but omitted in the response. 2. Format the provenance as follows: - ‘response: "[The entire response]".’ - ‘context: "[Relevant segment from the context that should have been matched]".’ Guidelines: - In all cases, ensure the provenance provided is clear, concise, and directly supports the selected "value". - The provenance should be an exact match or a direct comparison between the response and the context, following the formatting rules strictly.
@@
For these attributes, we ask the LLM to provide a value: True Positive (TP) - indicating accurate prediction, False Positive (FP) - indicating the detection of hallucination, or False Negative (FN) - indicating failed retrieval given the response and the evidence (from NVD+CWE+hyperlinks). Next, the LLM is asked to provide a rationale for the selected value, and finally the provenance attribute where the LLM is tasked to show evidence for its response.
@@
5 Evaluation of ProveRAG
Table 1: Comparison between Summarizing and Chunking techniques for retrieving the content for CVEs (Provenance Features: TP: accurate prediction, FP: detected hallucination, FN: failed retrieval)
@@ 5 Evaluation of ProveRAG
To answer each research question, we discuss the curation of the dataset for (RQ1) and provide results for different RAG techniques (RQ1 and RQ2). Furthermore, we support the efficacy of provenance by using evaluation metrics (RQ3), show how each URL attributed to the Gen. LLM’s response (RQ3), and finally we evaluate ProveRAG on another LLM (Llama-3.1) and we will show some specific examples of ProveRAG’s responses.
5.1 Dataset & Experiment Design
@@
5.2 Effectiveness of ProveRAG
5.2.1 Chunking vs. Summarizing Techniques
@@ 5.2 Effectiveness of ProveRAG
From the results, first we can see that for the prompt-only experiment, the performance is poor, for both exploitation and mitigation results. This shows the big limitation of LLMs’ knowledge with new, unseen data, as the entire dataset consists of 2024 vulnerabilities. This goes back to RQ1 and how we can deal with this problem. By integrating the retrieval techniques, we can see that the temporal alignment of the LLM is improved by the inclusion of the curated dataset, indicating that such an approach may alleviate limitations related to temporal alignment for cybersecurity applications.
For the exploitation results, all evaluated methodologies exhibited high accuracy, especially for NVD-Only which was limited to the information about the CVE’s exploitation on the NVD website. Another observation is that adding more data slightly decreases the chunking technique’s performance. In Contrast, for the more complex queries (i.e., mitigation information), significant variations are observed. The NVD-Only configuration achieves around 29% accuracy in providing mitigation strategies, which makes sense given that the NVD website lacks a dedicated mitigation section; yet, it does offer some recommendations for certain CVEs. When the CWE is added into the generation of results, we can see a big increase in TPs (from 139 to 262). However, adding additional URLs into the generation mix appears to diminish performance by about 7%, which suggests that an excess of links might degrade the quality of the retrieved data segments with the chunking technique. This indicates that providing excessive context can degrade performance when chunking is used.
On the other hand, the summarizing technique significantly outperforms the chunking technique when provided with the same set of links for data generation, showing an increase from 227 to 379 in TPs — a rise of over 30%. Therefore, these findings strongly support the conclusion that summarization is a better and more efficient technique. As a result, it has been incorporated into ProveRAG to help align the system to vulnerabilities discovered after training cut-off. Furthermore, this technique alleviates the context window problem (RQ2) for vulnerability analysis, as only the essential information (i.e., summaries) are put into the Gen. LLM’s prompt.
5.2.2 ProveRAG Results
@@
Table 2: ProveRAG results for the 482 CVEs
@@
5.3 Evaluation Metrics for Provenance
@@
5.3.1 Attribution:
To assess the attribution of each URL summary (i.e., NVD, CWE, and hyperlinks) and how much each contributed to the generation of the results, we made use of the Captum library [18]. This library offers implementations of various perturbation- and gradient-based attribution algorithms for PyTorch models. Attribution methods allow for model explainability by assigning scores to each input feature, representing its impact on the model’s output. In our experiment, we compute attribution scores using feature ablation, a perturbation-based method that evaluates the model on altered inputs. This method replaces each feature with its corresponding value from the baseline feature vector to assess its influence. Equation 1 assesses the contribution of different features to the final result.
$\Delta y=y(x)-y(x_{\setminus F})$ (1)
@@ 5.3.1 Attribution:
5.3.2 Rouge-L and Cosine Similarity:
@@
To evaluate the quality of the Provenance results, we incorporated Rouge-L and Cosine Similarity (based on embeddings) metrics into our analysis. Rouge-L focuses on the longest common subsequence, which measures the sequence similarity between the generated responses and the reference texts. On the other hand, Cosine Similarity leverages a Sentence Transformer model, specifically the multi-qa-mpnet-base-dot-v1, which is optimized for semantic search [30]. This model utilizes a 512-context window to compare the semantic congruence between the response and the evidence part of the provenance, ensuring that the evaluation captures not just the textual, but also the contextual and semantic accuracy of the information provided.
5.4 Provenance Quality
@@
To answer RQ3, we applied Rouge-L and Cosine Similarity to measure the quality of both the provided responses and their supporting provenance. These metrics help us assess the alignment between the system’s outputs and verifiable evidence when evaluating TPs, FPs, and FNs. Figure 3 illustrates these values. As anticipated, TPs consistently achieve the highest scores across both metrics, indicating greatest overlap between the generated response and the supporting evidence, which underscores the effectiveness of ProveRAG. Note that the scores of FNs is close to that of TPs. This is expected because the FNs often just omitted some details in the response when comparing it with the source context. There is no value for FN in the exploitation part, as the LLM made no FN predictions in that category. Detailed examples are provided in Section 5.7 to contextualize these values.
Figure 3: ProveRAG Exploitation & Mitigation Cosine Similarity and Rouge-L Values for #TPs, #FPs, and #FNs
@@
Furthermore, to analyze the effect of summary for each URL to the LLM’s response, we made use of the Captum library [18]. We calculate attribution by using the summaries from each relevant website as context for the generation prompt. If a website is deemed irrelevant, its attribution score is set to zero, as it does not provide usable information for the LLM. To facilitate comparison across CVEs, the attribution scores are normalized to unit norm. Figure 4 shows these attribution values. From the results, it is evident that for exploitation, the majority of information in the summaries originates from NVD. For mitigation summaries, most of CWE’s attention scores are higher than 0.9, whereas the values for NVD and the hyperlinks stay mostly unchanged. This indicates that the Gen. LLM put the most attention to the NVD summary for generating the exploitation response, and to the CWE summary for generating the mitigation strategies.
Figure 4: ProveRAG Attribution Results
5.5 Improvement w/ Structured Context
@@
In Sec. 5.2.2, we demonstrated that ProveRAG achieves 95% and 80% TPs (with NVD, CWE, and all hyperlinks in the NVD website). Therefore, we raise the question on whether this can be further improved. Note that CWE is addressing mitigation at the weakness level, not the vulnerability level. Through our research, we found Aqua Vulnerability Database [29], which gives a highly structured and detailed context for each CVE’s exploitation steps and mitigation strategies. Thus, we experimented ProveRAG by purposefully including Aqua in the retrieval summarizing process, and into the evaluation module for provenance. The results are shown in the bottom row of Table 2. The results demonstrate the significant increase in the exploitation and mitigation detection of ProveRAG-Aqua, with 99% and 97% for TPs, respectively. This shows that the Aqua database plays a pivotal role in enhancing ProveRAG’s performance by providing more detailed and structured context. These results are significantly better than the baseline experiments, where prompt-only achieved just 6% accuracy for mitigation strategies, and the retrieval chunking technique reached about 47%.
Table 3: Critical CVEs Relevancy with respect to the URLs
@@
Figure 5: ProveRAG-Aqua attribution results
@@
To further evaluate Aqua’s impact on the Gen. LLM’s responses, Figure 5 presents the attribution of each summary in ProveRAG-Aqua. The results reveal that, for the mitigation prompt, both Aqua and CWE exhibit a wide distribution of attention scores, with Aqua having a higher median around 0.7 compared to CWE’s median of 0.43. For the exploitation prompt, both Aqua and NVD display a broad range of attention scores, mostly positive, with Aqua again leading with a median near 0.7, while NVD’s median is around 0.3. This shows that when Aqua is included, it is the dominant source for the Gen. LLM for both exploitation and mitigation questions, with some contribution from NVD for exploitation and from CWE for mitigation strategies. Next, we will assess the performance of ProveRAG on another LLM.
5.6 ProveRAG’s Performance on Other LLMs (Generalizability)
To evaluate whether the proposed ProveRAG framework is generalizable across different models, we compare the gpt-4o-mini results with Llama-3.1-8B (both models have the same context window and knowledge cutoff before the year 2024). First, based on qualitative evaluation, we noticed that gpt-4o-mini model is significantly better in following instructions to generate the responses. Note that the size of gpt-4o-mini model is not disclosed, but we argue that it is a much richer model in following instructions and generating in the desired format. On the other hand, Llama-3.1-8B model tend to not pay as much attention on the guidelines in generating the outputs, and many times failing to output the value, rationale, and provenance attributes in the right format (such as providing multiple TP/FP/FNs for a single response and/or not showing evidence for provenance). Therefore, we added the following to the Evaluation Prompt of ProveRAG for the Llama-3.1-8B model to alleviate this issue:
Output in the following format: value: TP or FP or FN (ONLY one value for the entire response) rationale: [the rationale segment for the selected value] provenance: - response: [extract the relevant segment from the response segment]’ - context: [extract the relevant segment from the context segment]’
This indicates the limitation for LLMs with different sizes and capabilities, as they may struggle to follow all the instructions when tasked with a lengthy or complex set of instructions. However, through careful prompt engineering, this challenge can be substantially mitigated.
Another qualitative observation from Llama-3.1-8B is that the guardrails are more strict for this model. This was not the case for the gpt-4o-mini model as it never refused to generate the response for the Generation Prompt. However, Llama-3.1-8B generated the following sentence for all the CVEs for the prompt-only experiment:
I cannot provide information on how to exploit vulnerabilities. Is there anything else I can help you with?
This was mostly an issue for the prompt-only experiment, however, for ProveRAG (i.e., with retrieved and summarized information injected into the prompt), the Llama-3.1-8B only refused to generate the response for around 40 of the CVEs. Therefore, we did not include those 40 for provenance (as provenance requires a response to self-critique and provide evidence). This shows that different LLMs may have varying sets of guardrails, but with careful prompting, they can still benefit from the ProveRAG framework. Table 4 shows the results obtained. Note that the #GR indicates the number of guardrail responses, which means that the LLM refused to provide a response for them.
Table 4: Performance of ProveRAG for GPT-4o-mini vs. Llama3.1-8B Model
  Evidence Source (NVD + CWE + Relevant Refs.)
Model & Methodology   Exploitation  Mitigation
  #TPs #FPs #FNs  #TPs #FPs #FNs #GR
GPT-4o-mini
Prompt-only  51 (11%) 400 31    29 (6%) 335 118 0
ProveRAG  460 (95%) 22 0    384 (80%) 66 32 0
Llama3.1-8B
Prompt-only  0 0 0    0 0 0 482
ProveRAG  331 (69%) 73 38    357 (74%) 66 19 40
Based on the results, it is evident that ProveRAG performs significantly better with both the gpt-4o-mini and Llama-3.1-8B models compared to the baselines, with gpt-4o-mini showing particularly superior performance. This shows ProveRAG’s effectiveness in vulnerability analysis on another LLM, showcasing its ability to identify CVEs and conduct self-critique. Moreover, our findings indicated that Llama-3.1-8B tend to “forget” to adhere to all given instructions. Therefore, we recommend a more careful prompt engineering to enhance performance, which allows the model to better follow instructions and bypass overly rigid guardrails through the proposed ProveRAG framework. While the choice of model is flexible, we suggest using models with strong instruction-following capabilities for optimal outcomes.
Therefore, these findings support the generalizability of the proposed ProveRAG framework, emphasizing that with careful prompt engineering—taking into account the unique guardrails and instruction-following capabilities of each model—ProveRAG can effectively leverage different LLMs. To optimize the effectiveness of provenance, we recommend using models that excel at capturing complex instructions and understanding intricate token dependencies. In particular, prompts must be tailored to each model’s specific characteristics to fully leverage the benefits of ProveRAG. Next, we will show some specific examples on how the proposed ProveRAG framework works.
5.7 Specific Examples
In this section, we show how the evaluation module of ProveRAG works (RQ3). The first three examples present gpt-4o-mini responses with varying TP, FP, and FN predictions. Following these, we provide a comparative analysis of a specific CVE between the gpt-4o-mini and Llama-3.1-8B models. First, here is one example of gpt-4o-mini that shows a TP response: